HIGH RESOLUTION ELECTRONIC SPECTROSCOPY OF 9-FLUORENEMETHANOL IN THE GAS PHASE^{a}

\underline{DIANE M. MITCHELL}, DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, PA 15260; JAMES A.J. FITZPATRICK, Department of Chemistry, Carnegie Mellon University, 4400 Fifth Avenue, Pittsburgh, PA 15217.

High resolution fluorescence excitation spectra have been observed for the S_{1}-S_{0} transition of two conformers of 9fluorenemethanol. With the aid of theoretical calculations, the band at 33,788.8 cm^{−1} has been assigned to an asymmetric conformer while the two bands at 33,561.7 and 33,563.6 cm^{−1} have been assigned to the symmetric conformer. Possible explanations for the two symmetric bands will be explored. The discussion will include the feasibility of internal motion versus the prospect of a torsional motion(s) resulting in an asymmetric barrier.

aWork supported by NSF (CHE-0315584)